OpenStudy (anonymous):

Fred has been given the task of determining the height of a building. He walks 42m away from the base of the building and sites the top of the building at an angle of elevation of 59 degrees. Calculate the height of the building if Fred is 1.65m tall. Show all work.

OpenStudy (callisto):

 From the figure, h1 / 42 = tan 59 => h1 = 42tan59 m Height of the building = h1 +h2 = 42tan59 + 1.65 = Can you do it?